4(3x -2) > 12(x + 1)
Answer by DRichardson(63)

4(3x -2) > 12(x + 1)
Hello There,
With my beliefs this is how you would solve this problem.
Distribute:
4 (3x)- 4(2) > 12(x) + 12(1)
12x - 8 > 12x + 12

12x - 8 = 4x
12x + 12= 24x
4x>24x
Four x is less than twenty four x.
If you want to solve this for x = then you would do this:
24x ÷ 4x = 6x
x = 6
Hope this helps~!
Delilah R